### 15 Jan 2024 v5

# Reciprocity formulas for certain generalized Hardy-Berndt sums

Abstract.
In this paper, we establish some reciprocity formulas for certain generalized Hardy-Berndt sums by using the Fourier series technique and some properties of the periodic zeta function and the Lerch zeta function. It turns out that one of Hardy’s reciprocity theorems is deduced as a special case.
2020 Mathematics Subject Classification. 11B68; 11F20; 11M35.
Keywords. quasi-periodic Euler functions; Bernoulli functions; Hardy-Berndt sums; Hurwitz and Lerch zeta functions; Reciprocity formulas
1. Introduction
Let $\mathbb{N}$ be the set of positive integers, $\mathbb{N}_{0}$ the set of non-negative integers, $\mathbb{Z}$ the set of integers, $\mathbb{R}$ the set of real numbers, and $\mathbb{C}$ the set of complex numbers. Denote by $\{x\}$ the fractional part of $x\in\mathbb{R}$, and $((x))$ the sawtooth function (also called the first Bernoulli function) given by
$((x))=\begin{cases}\{x\}-\frac{1}{2},&\text{if $x\in\mathbb{R}\setminus\mathbb{Z}$},\\ 0,&\text{if $x\in\mathbb{Z}$}.\end{cases}$
The classical Dedekind sums $s(a,b)$, arising in the transformation formulas of the logarithm of the Dedekind eta function, is defined for $a,b\in\mathbb{Z}$ with $b\not=0$ by
$s(a,b)=\sum_{r=0}^{|b|-1}\biggl{(}\biggl{(}\frac{r}{b}\biggl{)}\biggl{)}\biggl{(}\biggl{(}\frac{ar}{b}\biggl{)}\biggl{)}.$ (1.1)
It is well known that one of the most intriguing and important features for the sums (1.1) is the following Dedekind’s [12] reciprocity theorem. If $a$ and $b$ are two relatively prime positive integers, then
$s(a,b)+s(b,a)=-\frac{1}{4}+\frac{1}{12}\biggl{(}\frac{a}{b}+\frac{b}{a}+\frac{1}{ab}\biggl{)}.$ (1.2)
In the year 1905, Hardy [14] first provided a different proof of the reciprocity formula (1.2) which does not depend on the theory of the Dedekind eta function, and stated some reciprocity theorems for some similar arithmetic sums to the sums (1.1). For example, Hardy [14, Equation (ix)] shown that for two relatively prime odd positive integers $a,b$,
$s_{5}(a,b)+s_{5}(b,a)=\frac{1}{2}-\frac{1}{2ab},$ (1.3)
where $s_{5}(a,b)$ is the Hardy sums (also known the Hardy-Berndt sums) given for $a,b\in\mathbb{Z}$ with $b\not=0$ by
$s_{5}(a,b)=\sum_{r=0}^{|b|-1}(-1)^{r+[\frac{ar}{b}]}\biggl{(}\biggl{(}\frac{r}{b}\biggl{)}\biggl{)},$ (1.4)
in which $[x]$ is the floor function satisfying $[x]=x-\{x\}$ for $x\in\mathbb{R}$. Just as Dedekind [12] deduced the reciprocity formula (1.2) from his transformation formula of the logarithm of the Dedekind eta function, Berndt [5] in 1978 used his transformation formulas of the logarithms of the classical theta-functions to prove three of Hardy’s reciprocity theorems. Goldberg [13] in 1981 used Berndt’s [5] transformation formulas to prove the reciprocity formula (1.3). In fact, Goldberg [13] also presented some three-term and mixed three-term reciprocity formulas for those sums considered by Hardy [14] and Berndt [5]. For some different proofs and extensions of Hardy’s reciprocity theorems, one is referred to [6, 7, 8, 11, 20, 23].
The object of the present paper is to give some new extensions of the reciprocity formula (1.3). To do so, we consider the following generalized Hardy-Berndt sums
$S_{m,n}^{(5)}\left(\begin{matrix}a&b&c\\ x&y&z\end{matrix}\right)=\sum_{r=0}^{|c|-1}(-1)^{r}\overline{E}_{m}\biggl{(}\frac{a(r+z)}{c}-x\biggl{)}\overline{B}_{n}\biggl{(}\frac{b(r+z)}{c}-y\biggl{)},$ (1.5)
where $m,n\in\mathbb{N}_{0}$, $a,b,c\in\mathbb{Z}$, $x,y,z\in\mathbb{R}$ with $c\not=0$, $\overline{E}_{n}(x)$ is the $n$-th quasi-periodic Euler function given for $n\in\mathbb{N}_{0}$, $x\in\mathbb{R}$ by
$\overline{E}_{0}(x)=\begin{cases}(-1)^{[x]},&\text{if $x\in\mathbb{R}\setminus\mathbb{Z}$},\\ 0,&\text{if $x\in\mathbb{Z}$},\end{cases}\quad\overline{E}_{n}(x)=(-1)^{[x]}E_{n}(\{x\})\quad(n\geq 1),$ (1.6)
$\overline{B}_{n}(x)$ is the $n$-th Bernoulli function given for $n\in\mathbb{N}_{0}$, $x\in\mathbb{R}$ by
$\overline{B}_{0}(x)=1,\quad\overline{B}_{1}(x)=((x)),\quad\overline{B}_{n}(x)=B_{n}(\{x\})\quad(n\geq 2),$ (1.7)
in which $E_{n}(x)$ is the Euler polynomial of degree $n$, $B_{n}(x)$ is the Bernoulli polynomial of degree $n$. By using the Fourier series technique and some properties of the periodic zeta function and the Lerch zeta function, we establish some reciprocity formulas for the generalized Hardy-Berndt sums (1.5). As immediate consequences of the results presented here, we rediscover the reciprocity formula (1.3).
We now state our first result as follows.
Theorem 1.1.
Let $m,n\in\mathbb{N}_{0}$, $a,b\in\mathbb{Z}$ with $a\equiv b\equiv 1$ (mod $2$). Then, for $x,y,z\in\mathbb{R}$,
$2b^{n}\mathrm{sgn}(b)\sum_{j=0}^{m}\binom{m}{j}\frac{(-1)^{j}a^{m-j}}{m+n+1-j}S_{j,m+n+1-j}^{(5)}\left(\begin{matrix}a&1&b\\ y&x&z\end{matrix}\right)$ (1.10)
$\quad+2a^{m}\mathrm{sgn}(a)\sum_{j=0}^{n}\binom{n}{j}\frac{(-1)^{j}b^{n-j}}{m+n+1-j}S_{j,m+n+1-j}^{(5)}\left(\begin{matrix}b&1&a\\ z&x&y\end{matrix}\right)$ (1.13)
$=(-1)^{m+n}\overline{E}_{m}(ax-y)\overline{E}_{n}(bx-z)$
$\quad+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)(-1)^{ax+bx-y-z}\delta_{\mathbb{Z}}(ax-y)\delta_{\mathbb{Z}}(bx-z)$
$\quad+\frac{(-1)^{m}2m!n!(a,b)^{m+n+2}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{az-by}{(a,b)}\biggl{)},$ (1.14)
where $\mathrm{sgn}(x)$ indicates the sign of $x$ given for $x\in\mathbb{R}\setminus\{0\}$ by $\mathrm{sgn}(x)=x/|x|$, $\delta_{l,k}$ is the Kronecker delta function given for $l,k\in\mathbb{N}_{0}$ by $\delta_{l,k}=1$ or 00 according to $l=k$ or $l\not=k$, $\delta_{\mathbb{Z}}(x)=1$ or 00 according to $x\in\mathbb{Z}$ or $x\not\in\mathbb{Z}$, $(a,b)$ denotes the greatest common divisor of two integers $a,b$.
In particular, taking $m=n=x=y=z=0$ in Theorem 1.1, we can remove the hypothesis $(a,b)=1$ in the reciprocity formula (1.3), and obtain the following result.
Corollary 1.2.
Let $a,b\in\mathbb{N}$ with $a\equiv b\equiv 1$ (mod $2$). Then
$s_{5}(a,b)+s_{5}(b,a)=\frac{1}{2}-\frac{(a,b)^{2}}{2ab}.$ (1.15)
More generally, we have the following reciprocity formula.
Theorem 1.3.
Let $m,n\in\mathbb{N}_{0}$, $a,b,c\in\mathbb{Z}$ with $a\equiv b\equiv 1$ (mod $2$) and $c\not=0$. Then, for $x,y,z\in\mathbb{R}$,
$2b^{n}\mathrm{sgn}(bc)\sum_{j=0}^{m}\binom{m}{j}\frac{(-1)^{j}a^{m-j}}{c^{m+n-j}(m+n+1-j)}S_{j,m+n+1-j}^{(5)}\left(\begin{matrix}a&c&b\\ y&x&z\end{matrix}\right)$ (1.18)
$\quad+2a^{m}\mathrm{sgn}(ac)\sum_{j=0}^{n}\binom{n}{j}\frac{(-1)^{j}b^{n-j}}{c^{m+n-j}(m+n+1-j)}S_{j,m+n+1-j}^{(5)}\left(\begin{matrix}b&c&a\\ z&x&y\end{matrix}\right)$ (1.21)
$=(-1)^{m+n}\sum_{r=0}^{|c|-1}\overline{E}_{m}\biggl{(}-\frac{a(r-x)}{c}-y\biggl{)}\overline{E}_{n}\biggl{(}-\frac{b(r-x)}{c}-z\biggl{)}$
$\quad+\frac{(-1)^{m}2\mathrm{sgn}(c)m!n!(a,b)^{m+n+2}c}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{az-by}{(a,b)}\biggl{)}$
$\quad+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)\sum_{r=0}^{|c|-1}(-1)^{\frac{(a+b)(r-x)}{c}+y+z}$
×$\qquad\times\delta_{\mathbb{Z}}\biggl{(}\frac{a(r-x)}{c}+y\biggl{)}\delta_{\mathbb{Z}}\biggl{(}\frac{b(r-x)}{c}+z\biggl{)},$ (1.22)
where $\mathrm{sgn}(x)$, $(a,b)$, $\delta_{l,k}$, $\delta_{\mathbb{Z}}(x)$ are as in (1.10).
It becomes obvious that the case $m=n=x=y=z=0$ in Theorem 1.3 gives another extension of the reciprocity formula (1.3) as follows.
Corollary 1.4.
Let $a,b,c\in\mathbb{N}$ with $a\equiv b\equiv 1$ (mod $2$) and $(a,b)=(b,c)=(a,c)=1$. Then
$s_{5}(a,c,b)+s_{5}(b,c,a)=\frac{1}{2}\sum_{r=1}^{c-1}(-1)^{[\frac{ar}{c}]+[\frac{br}{c}]}-\frac{c}{2ab}+\frac{1}{2},$ (1.23)
where $s_{5}(a,b,c)$ is given for $a,b,c\in\mathbb{Z}$ with $c\not=0$ by
$s_{5}(a,b,c)=\sum_{r=0}^{|c|-1}(-1)^{r+[\frac{ar}{c}]}\biggl{(}\biggl{(}\frac{br}{c}\biggl{)}\biggl{)},$ (1.24)
the sum on the right hand side of (1.23) vanishes when $c=1$.
It should be noted that replacing $a$ by $c$ and $c$ by $a$ in Corollary 1.4, and then using the relations shown in [20, Equations (3.7), (3.8) and (3.9)], we can reobtain Goldberg’s [13, Theorem 5.6] mixed three-term reciprocity formula.
This paper is organized as follows. In the second section, we give some auxiliary results. The third section concentrates on the feature that has contributed to the detailed proofs of Theorems 1.1 and 1.3.
2. Some auxiliary results
Before giving the proofs of Theorems 1.1 and 1.3, we need the following auxiliary lemmas. For convenience, in the following we always denote by $\mathrm{i}$ the square root of $-1$ such that $\mathrm{i}^{2}=-1$, $\Gamma(s)$ the gamma function defined on $s\in\mathbb{C}$. For the sake of convergence, the sum
$\sum_{d=-\infty}^{+\infty}\frac{1}{d+a}\quad(a\not\in\mathbb{Z})$
is interpreted as
$\lim_{N\rightarrow\infty}\sum_{d=-N}^{N}\frac{1}{d+a}.$
This ensures that the $n$-th quasi-periodic Euler function $\overline{E}_{n}(x)$ defined in (1.6) can be given for $n\in\mathbb{N}_{0}$, $x\in\mathbb{R}$ by the Fourier series (see, e.g., [3, Theorem 1.2] or [19, Theorem 2.2])
$\overline{E}_{n}(x)=\frac{2n!}{(2\pi\mathrm{i})^{n+1}}\sum_{k=-\infty}^{+\infty}\frac{e^{2\pi\mathrm{i}(k-\frac{1}{2})x}}{\bigl{(}k-\frac{1}{2}\bigl{)}^{n+1}},$ (2.1)
the $n$-th Bernoulli function $\overline{B}_{n}(x)$ defined in (1.7) can be given for $n\in\mathbb{N}$, $x\in\mathbb{R}$ by the Fourier series (see, e.g., [2, Theorem 12.19] or [3, Theorem 1.1] or [19, Theorem 2.1])
$\overline{B}_{n}(x)=-\frac{n!}{(2\pi\mathrm{i})^{n}}\sideset{}{{}^{\prime}}{\sum}_{k=-\infty}^{+\infty}\frac{e^{2\pi\mathrm{i}kx}}{k^{n}},$ (2.2)
where the dash denotes throughout that undefined terms are excluded from the sum. We now present the following results.
Lemma 2.1.
(Parseval’s formula) Suppose that $F(\theta)$ and $G(\theta)$ are two Riemann integrable, complex-valued functions on $\mathbb{R}$ of period $2\pi$ with the Fourier series
$F(\theta)=\sum_{n=-\infty}^{+\infty}a_{n}e^{\mathrm{i}n\theta},$
and
$G(\theta)=\sum_{n=-\infty}^{+\infty}b_{n}e^{\mathrm{i}n\theta}.$
Then
[…]
where the horizontal bars indicate complex conjugation.
Proof.
See [10, Proposition 3.1.10] or [24, p. 81] for details. ∎
Lemma 2.2.
(Raabe’s multiplication formula) Let $n,a\in\mathbb{N}$. Then, for $x\in\mathbb{R}$,
$a^{n-1}\sum_{r=0}^{a-1}\overline{B}_{n}\biggl{(}x+\frac{r}{a}\biggl{)}=\overline{B}_{n}(ax).$ (2.4)
Proof.
See [21] or [16, Equation (28)] for details. ∎
Lemma 2.3.
Let $m,n\in\mathbb{N}$, $d,x,y\in\mathbb{R}$ with $d\not=x\not=y$. Then
$\frac{1}{(d-x)^{m}(d-y)^{n}}$ $=$ $\sum_{j=1}^{m}\binom{m+n-j-1}{n-1}\frac{(-1)^{m-j}}{(x-y)^{m+n-j}(d-x)^{j}}$ (2.5)
$+\sum_{j=1}^{n}\binom{m+n-j-1}{m-1}\frac{(-1)^{n-j}}{(y-x)^{m+n-j}(d-y)^{j}}.$
Proof.
See [17, Lemma 2.2] for details. ∎
Lemma 2.4.
Let $q,j\in\mathbb{N}$ with $q\geq 2$, and let $\theta_{r}$ be a real-valued function defined on $r\in\mathbb{N}$ such that $\theta_{r}\not=0,\pm q,\pm 2q,\ldots$. Then
$\frac{\partial^{j-1}}{\partial a^{j-1}}\bigl{(}\cot(\pi a)\bigl{)}\biggl{|}_{a=\frac{\theta_{r}}{q}}=\frac{\delta_{1,j}}{\mathrm{i}}+2^{j}\pi^{j-1}\mathrm{i}^{j-2}F\biggl{(}\frac{\theta_{r}}{q},1-j\biggl{)},$ (2.6)
where $F(x,s)$ is the periodic zeta function given for $x\in\mathbb{R}$, $s\in\mathbb{C}$ by (see, e.g., [2, p. 257])
$F(x,s)=\sum_{n=1}^{\infty}\frac{e^{2\pi\mathrm{i}nx}}{n^{s}}\quad(\Re(s)>1).$
Proof.
See [15, Equation (2.28)] for details. ∎
Lemma 2.5.
Let $j\in\mathbb{N}$, $b,r\in\mathbb{Z}$ with $b\not=0$. Then
$\sum_{d=-\infty}^{+\infty}\frac{1}{\bigl{(}d+\frac{r}{b}-\frac{1}{2b}\bigl{)}^{j}}=\frac{(2\pi\mathrm{i})^{j}\mathrm{sgn}(b)}{2(j-1)!b^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}\overline{E}_{j-1}\biggl{(}\frac{l}{b}\biggl{)}.$ (2.7)
Proof.
Since $\cot(a)$ has the following expression in partial fractions (see, e.g., [1, p. 75] or [22, p. 327])
$\cot(a)$ $=$ $\frac{1}{a}+2a\sum_{d=1}^{\infty}\frac{1}{a^{2}-d^{2}\pi^{2}}$
$=$ $\sum_{d=-\infty}^{+\infty}\frac{1}{a+d\pi}\quad(a\not=0,\pm\pi,\pm 2\pi,\ldots),$
we know that for $j\in\mathbb{N}$, $b,r\in\mathbb{Z}$ with $b\not=0$,
$\frac{\partial^{j-1}}{\partial a^{j-1}}\bigl{(}\cot(\pi a)\bigl{)}\biggl{|}_{a=\frac{r}{b}-\frac{1}{2b}}=\frac{(-1)^{j-1}(j-1)!}{\pi}\sum_{d=-\infty}^{+\infty}\frac{1}{\bigl{(}d+\frac{r}{b}-\frac{1}{2b}\bigl{)}^{j}}.$ (2.8)
By taking $\theta_{r}=q(\frac{r}{b}-\frac{1}{2b})$ in Lemma 2.4, we conclude from (2.8) that
$\sum_{d=-\infty}^{+\infty}\frac{1}{\bigl{(}d+\frac{r}{b}-\frac{1}{2b}\bigl{)}^{j}}=-\delta_{1,j}\pi\mathrm{i}+\frac{(-1)^{j}(2\pi\mathrm{i})^{j}}{(j-1)!}F\biggl{(}\frac{r}{b}-\frac{1}{2b},1-j\biggl{)}.$ (2.9)
It is clear that for $j\in\mathbb{N}_{0}$ (see, e.g., [9, Corollary 3]),
$\eta(-j,x)=\frac{1}{2}E_{j}(x),$ (2.10)
where $\eta(s,x)$ is the alternating Hurwitz zeta function given for $s\in\mathbb{C}$, $x\in\mathbb{R}$ with $x>0$ by
$\eta(s,x)=\sum_{n=0}^{\infty}(-1)^{n}\frac{1}{(n+x)^{s}}\quad(\Re(s)>0).$
Hence, using the familiar division algorithm stated in [2, Theorem 1.14] and (2.10), the periodic zeta function on the right hand side of (2.9) can be rewritten as
$F\biggl{(}\frac{r}{b}-\frac{1}{2b},1-j\biggl{)}$ $=$ $\sum_{l=1}^{|b|}\sum_{m=0}^{\infty}\frac{e^{\frac{2\pi\mathrm{i}(m|b|+l)r}{b}}e^{-\frac{\pi\mathrm{i}(m|b|+l)}{b}}}{(m|b|+l)^{1-j}}$ (2.11)
$=$ $\sum_{l=1}^{|b|}e^{\frac{2\pi\mathrm{i}lr}{b}-\frac{\pi\mathrm{i}l}{b}}\sum_{m=0}^{\infty}(-1)^{m}\frac{1}{(m|b|+l)^{1-j}}$
$=$ $\frac{1}{2|b|^{1-j}}\sum_{l=1}^{|b|}e^{\frac{2\pi\mathrm{i}lr}{b}-\frac{\pi\mathrm{i}l}{b}}E_{j-1}\biggl{(}\frac{l}{|b|}\biggl{)}$
$=$ $-\frac{1}{2|b|^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}E_{j-1}\biggl{(}1-\frac{l}{|b|}\biggl{)}.$
Since the Euler polynomials satisfy the difference equation (see, e.g., [1, p. 804])
$E_{n}(1-x)=(-1)^{n}E_{n}(x)\quad(n\in\mathbb{N}_{0}),$
we see from $E_{0}(x)=1$ that for $j\in\mathbb{N}$, $l\in\mathbb{N}_{0}$, $b\in\mathbb{Z}$ with $10\leq l\leq|b|-$ 1,
$E_{j-1}\biggl{(}1-\frac{l}{|b|}\biggl{)}=(-1)^{j-1}\overline{E}_{j-1}\biggl{(}\frac{l}{|b|}\biggl{)}+\delta_{l,0}\delta_{1,j}.$ (2.12)
It follows from (2.11) and (2.12) that
$F\biggl{(}\frac{r}{b}-\frac{1}{2b},1-j\biggl{)}=\frac{(-1)^{j}}{2|b|^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}\overline{E}_{j-1}\biggl{(}\frac{l}{|b|}\biggl{)}-\frac{1}{2}\delta_{1,j}.$ (2.13)
It is easily seen from (2.1) that for $n\in\mathbb{N}_{0}$, $x\in\mathbb{R}$,
$\overline{E}_{n}(-x)=\begin{cases}\overline{E}_{n}(x),&x\in\mathbb{Z},\\ (-1)^{n+1}\overline{E}_{n}(x),&x\not\in\mathbb{Z}.\end{cases}$ (2.14)
With the help of (2.14), we rewrite (2.13) as
$F\biggl{(}\frac{r}{b}-\frac{1}{2b},1-j\biggl{)}$
$=\frac{(-1)^{j}\mathrm{sgn}(b)}{2b^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}\overline{E}_{j-1}\biggl{(}\frac{l}{b}\biggl{)}$
$\quad+\frac{(-1)^{j}(\mathrm{sgn}(b))^{j-1}-\mathrm{sgn}(b)}{2b^{1-j}}\bigl{(}E_{j-1}(0)-\delta_{1,j}\bigl{)}-\frac{1}{2}\delta_{1,j}.$ (2.15)
Note that for $j\in\mathbb{N}$, (see, e.g., [1, p. 805])
$E_{j-1}(0)=-\frac{2(2^{j}-1)}{j}B_{j},$ (2.16)
and
$B_{2j+1}=0,$ (2.17)
where $B_{j}$ is the $j$-th Bernoulli number. Therefore, by applying (2.16) and (2.17) to the right hand side of (2), we have
$F\biggl{(}\frac{r}{b}-\frac{1}{2b},1-j\biggl{)}=\frac{(-1)^{j}\mathrm{sgn}(b)}{2b^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}\overline{E}_{j-1}\biggl{(}\frac{l}{b}\biggl{)}-\frac{1}{2}\delta_{1,j},$
from which and (2.9) we obtain (2.7) immediately. This completes the proof of Lemma 2.5. ∎
Lemma 2.6.
(Lerch’s functional equation) If $10<x<$ 1 and $0<a\leq 1$ then for all $s\in\mathbb{C}$,
$\phi(x,a,1-s)=\frac{\Gamma(s)}{(2\pi)^{s}}\bigl{(}e^{\frac{\pi\mathrm{i}s}{2}-2\pi\mathrm{i}ax}\phi(-a,x,s)+e^{-\frac{\pi\mathrm{i}s}{2}+2\pi\mathrm{i}a(1-x)}\phi(a,1-x,s)\bigl{)},$ (2.18)
where $\phi(x,a,s)$ is the Lerch zeta function given for $x,a\in\mathbb{R}$, $s\in\mathbb{C}$ with $0<a\leq 1$ by
$\phi(x,a,s)=\sum_{n=0}^{\infty}\frac{e^{2\pi\mathrm{i}nx}}{(n+a)^{s}}\quad(\Re(s)>1).$
Proof.
See [4, 18] for details. ∎
Lemma 2.7.
Let $j\in\mathbb{N}$, $b,r\in\mathbb{Z}$ with $b\not=0$. Then, for $x\in\mathbb{R}$,
$\sum_{d=-\infty}^{+\infty}\frac{e^{2\pi\mathrm{i}dx}}{\bigl{(}d+\frac{r}{b}-\frac{1}{2b}\bigl{)}^{j}}$
$=\frac{(2\pi\mathrm{i})^{j}\mathrm{sgn}(b)}{2(j-1)!b^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}(l+x)r}{b}+\frac{\pi\mathrm{i}(l+x)}{b}}\overline{E}_{j-1}\biggl{(}\frac{l+x}{b}\biggl{)}.$ (2.19)
Proof.
We first consider the case $x\in\mathbb{Z}$. Since $l+x$ runs over a complete residue system modulo $|b|$ as $l$ does. Hence, write
$l+x=q_{l}|b|+k_{r},$
where $q_{l}\in\mathbb{Z}$, $k_{r}\in\mathbb{N}_{0}$ with $10\leq k_{r}\leq|b|-$ 1, we know from (2.1) and Lemma 2.5 that Lemma 2.7 holds true in the case when $x\in\mathbb{Z}$. We next discuss the case $x\not\in\mathbb{Z}$. It suffices to prove that Lemma 2.7 holds true in the case when $10<x<$ 1. Obviously, in this case, we have
$\sum_{d=-\infty}^{+\infty}\frac{e^{2\pi\mathrm{i}dx}}{\bigl{(}d+\frac{r}{b}-\frac{1}{2b}\bigl{)}^{j}}$
$=e^{-2\pi\mathrm{i}x[\frac{r}{b}-\frac{1}{2b}]}\biggl{(}\sum_{d=0}^{\infty}\frac{e^{2\pi\mathrm{i}dx}}{(d+\{\frac{r}{b}-\frac{1}{2b}\})^{j}}+(-1)^{j}e^{-2\pi\mathrm{i}x}\sum_{d=0}^{\infty}\frac{e^{-2\pi\mathrm{i}dx}}{(d+1-\{\frac{r}{b}-\frac{1}{2b}\})^{j}}\biggl{)}$
$=e^{-2\pi\mathrm{i}x[\frac{r}{b}-\frac{1}{2b}]}$
×$\quad\times\biggl{(}\phi\biggl{(}x,\biggl{\{}\frac{r}{b}-\frac{1}{2b}\biggl{\}},j\biggl{)}+(-1)^{j}e^{-2\pi\mathrm{i}x}\phi\biggl{(}-x,1-\biggl{\{}\frac{r}{b}-\frac{1}{2b}\biggl{\}},j\biggl{)}\biggl{)},$ (2.20)
where $\phi(x,a,s)$ is as in (2.18). If we take $x=1-\{\frac{r}{b}-\frac{1}{2b}\}$ and $s=j$ and replace $a$ by $x$ in Lemma 2.6, then we have
$\phi\biggl{(}1-\biggl{\{}\frac{r}{b}-\frac{1}{2b}\biggl{\}},x,1-j\biggl{)}$
$=\frac{(j-1)!}{(2\pi)^{j}}\biggl{(}e^{\frac{\pi\mathrm{i}j}{2}-2\pi\mathrm{i}x(1-\{\frac{r}{b}-\frac{1}{2b}\})}\phi\biggl{(}-x,1-\biggl{\{}\frac{r}{b}-\frac{1}{2b}\biggl{\}},j\biggl{)}$
$\quad+e^{-\frac{\pi\mathrm{i}j}{2}+2\pi\mathrm{i}x\{\frac{r}{b}-\frac{1}{2b}\}}\phi\biggl{(}x,\biggl{\{}\frac{r}{b}-\frac{1}{2b}\biggl{\}},j\biggl{)}\biggl{)}.$ (2.21)
Multiplying both sides of (2) by $e^{\frac{\pi\mathrm{i}j}{2}-2\pi\mathrm{i}x\{\frac{r}{b}-\frac{1}{2b}\}}$, and it follows from (2) that
$\sum_{d=-\infty}^{+\infty}\frac{e^{2\pi\mathrm{i}dx}}{\bigl{(}d+\frac{r}{b}-\frac{1}{2b}\bigl{)}^{j}}=e^{-\frac{2\pi\mathrm{i}rx}{b}+\frac{\pi\mathrm{i}x}{b}}\frac{(2\pi\mathrm{i})^{j}\phi(1-\{\frac{r}{b}-\frac{1}{2b}\},x,1-j)}{(j-1)!}.$ (2.22)
Note that from the division algorithm, (2.10) and (2.14) we have
$\phi\biggl{(}1-\biggl{\{}\frac{r}{b}-\frac{1}{2b}\biggl{\}},x,1-j\biggl{)}$
$=\sum_{l=0}^{|b|-1}\sum_{m=0}^{\infty}\frac{e^{-\frac{2\pi\mathrm{i}(m|b|+l)r}{b}+\frac{\pi\mathrm{i}(m|b|+l)}{b}}}{(m|b|+l+x)^{1-j}}$
$=\frac{1}{2|b|^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}E_{j-1}\biggl{(}\frac{l+x}{|b|}\biggl{)}$
$=\frac{\mathrm{sgn}(b)}{2b^{1-j}}\sum_{l=0}^{|b|-1}e^{-\frac{2\pi\mathrm{i}lr}{b}+\frac{\pi\mathrm{i}l}{b}}\overline{E}_{j-1}\biggl{(}\frac{l+x}{b}\biggl{)}.$ (2.23)
Therefore, inserting (2) into (2.22), we say that Lemma 2.7 holds true in the case when $x\not\in\mathbb{Z}$. This completes the proof of Lemma 2.7. ∎
3. The proofs of Theorems 1.1 and 1.3
The proof of Theorem 1.1. Let
$f(x)=e^{\pi\mathrm{i}(ax+bx+y+z)}\overline{E}_{m}(ax+y)\overline{E}_{n}(bx+z).$
Since $f(x)$ is of bounded variation on every finite interval, $f(x)$ may be expanded in a Fourier series
$\frac{f(x^{+})+f(x^{-})}{2}=\sum_{k=-\infty}^{+\infty}c_{k}(m,n|a,b,y,z)e^{2\pi\mathrm{i}kx},$ (3.1)
where the Fourier coefficients $c_{k}(m,n|a,b,y,z)$ are determined by
[…]
If we make the change of the variable $\theta=2\pi x$ in Lemma 2.1 and set
$F(2\pi x)=e^{\pi\mathrm{i}(ax+y)}\overline{E}_{m}(ax+y),\quad\overline{G(2\pi x)}=e^{\pi\mathrm{i}(bx+z)}\overline{E}_{n}(bx+z)e^{-2\pi\mathrm{i}kx},$
then we discover from (2.3) and (3.2) that
$c_{k}(m,n|a,b,y,z)=\frac{4m!n!}{(2\pi\mathrm{i})^{m+n+2}}\underset{al+bj=k}{\sum_{l=-\infty}^{+\infty}\sum_{j=-\infty}^{+\infty}}\frac{e^{2\pi\mathrm{i}(ly+jz)}}{(l-\frac{1}{2})^{m+1}(j-\frac{1}{2})^{n+1}}.$ (3.3)
It is well known that the Diophantine equation $ax+by=k$ is solvable if and only if $(a,b)\mid k$, and in the case when it is solvable, all solutions of $ax+by=k$ are given by
$x=\frac{k\overline{a}}{(a,b)}+\frac{bd}{(a,b)},\quad y=\frac{k\overline{b}}{(a,b)}-\frac{ad}{(a,b)},$
where $\overline{a},\overline{b},d\in\mathbb{Z}$ satisfy $a\overline{a}+b\overline{b}=(a,b)$. It follows from (3.3) that if $(a,b)\nmid k$ then
$c_{k}(m,n|a,b,y,z)=0,$ (3.4)
and if $(a,b)\mid k$ then
$c_{k}(m,n|a,b,y,z)$ (3.5)
$=\frac{4m!n!(a,b)^{m+n+2}e^{\frac{2\pi\mathrm{i}k(\overline{a}y+\overline{b}z)}{(a,b)}}}{(2\pi\mathrm{i})^{m+n+2}}$
×$\quad\times\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}d(by-az)}{(a,b)}}}{\bigl{(}k\overline{a}+bd-\frac{(a,b)}{2}\bigl{)}^{m+1}\bigl{(}k\overline{b}-ad-\frac{(a,b)}{2}\bigl{)}^{n+1}}$
$=\frac{(-1)^{n+1}4m!n!(a,b)^{m+n+2}e^{\frac{2\pi\mathrm{i}k(\overline{a}y+\overline{b}z)}{(a,b)}}}{a^{n+1}b^{m+1}(2\pi\mathrm{i})^{m+n+2}}$
×$\quad\times\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}d(by-az)}{(a,b)}}}{\bigl{(}d+\frac{k\overline{a}}{b}-\frac{(a,b)}{2b}\bigl{)}^{m+1}\bigl{(}d-\frac{k\overline{b}}{a}+\frac{(a,b)}{2a}\bigl{)}^{n+1}}.$
In particular, since $\overline{a}$ and $\overline{b}$ have the opposite parity in the case when $a\equiv b\equiv 1$ (mod $2$), by setting $k=\frac{a+b}{2}$ in (3.5), in view of (2.1), we have
$c_{\frac{a+b}{2}}(m,n|a,b,y,z)$
$=\frac{(-1)^{n+1}4m!n!(a,b)^{m+n+2}e^{\frac{\pi\mathrm{i}(a+b)(\overline{a}y+\overline{b}z)}{(a,b)}}}{a^{n+1}b^{m+1}(2\pi\mathrm{i})^{m+n+2}}\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}d(by-az)}{(a,b)}}}{\bigl{(}d+\frac{\overline{a}-\overline{b}}{2}\bigl{)}^{m+n+2}}$
$=\frac{(-1)^{n+1}4m!n!(a,b)^{m+n+2}e^{\frac{\pi\mathrm{i}(a+b)(\overline{a}y+\overline{b}z)}{(a,b)}}}{a^{n+1}b^{m+1}(2\pi\mathrm{i})^{m+n+2}}\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}d(by-az)}{(a,b)}}}{\bigl{(}d+\frac{\overline{a}-\overline{b}+1}{2}-\frac{1}{2}\bigl{)}^{m+n+2}}$
$=\frac{(-1)^{n+1}4m!n!(a,b)^{m+n+2}e^{\pi\mathrm{i}(y+z)}}{a^{n+1}b^{m+1}(2\pi\mathrm{i})^{m+n+2}}\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}(d-\frac{1}{2})(by-az)}{(a,b)}}}{\bigl{(}d-\frac{1}{2}\bigl{)}^{m+n+2}}$
$=\frac{(-1)^{n+1}2m!n!(a,b)^{m+n+2}e^{\pi\mathrm{i}(y+z)}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{by-az}{(a,b)}\biggl{)}.$ (3.6)
We next consider the case $k\not=\frac{a+b}{2}$ in (3.5). By taking
$x=-\frac{k\overline{a}}{b}+\frac{(a,b)}{2b},\quad y=\frac{k\overline{b}}{a}-\frac{(a,b)}{2a}$
in Lemma 2.2, we obtain from (3.5) that for $(a,b)\mid k$ with $k\not=\frac{a+b}{2}$,
$c_{k}(m,n|a,b,y,z)$
$=\frac{(-1)^{n+1}4m!n!(a,b)^{m+n+2}e^{\frac{2\pi\mathrm{i}k(\overline{a}y+\overline{b}z)}{(a,b)}}}{a^{n+1}b^{m+1}(2\pi\mathrm{i})^{m+n+2}}$
×$\quad\times\biggl{(}\sum_{j=1}^{m+1}\binom{m+n+1-j}{n}(-1)^{m+1-j}\biggl{(}-\frac{ab}{k(a,b)-(a,b)\frac{a+b}{2}}\biggl{)}^{m+n+2-j}$
×$\qquad\quad\times\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}d(by-az)}{(a,b)}}}{\bigl{(}d+\frac{k\overline{a}}{b}-\frac{(a,b)}{2b}\bigl{)}^{j}}$
$\qquad+\sum_{j=1}^{n+1}\binom{m+n+1-j}{m}(-1)^{n+1-j}\biggl{(}\frac{ab}{k(a,b)-(a,b)\frac{a+b}{2}}\biggl{)}^{m+n+2-j}$
×$\qquad\quad\times\sum_{d=-\infty}^{+\infty}\frac{e^{\frac{2\pi\mathrm{i}d(by-az)}{(a,b)}}}{\bigl{(}d-\frac{k\overline{b}}{a}+\frac{(a,b)}{2a}\bigl{)}^{j}}\biggl{)}.$ (3.7)
With the help of Lemma 2.7, we discover that (3) can be rewritten in the following way
$c_{k}(m,n|a,b,y,z)$
$=2b^{n}m!n!(a,b)\mathrm{sgn}(b)\sum_{j=1}^{m+1}\binom{m+n+1-j}{n}\frac{a^{m+1-j}}{(j-1)!\bigl{(}2\pi\mathrm{i}(k-\frac{a+b}{2})\bigl{)}^{m+n+2-j}}$
×$\qquad\times\sum_{l=0}^{|\frac{b}{(a,b)}|-1}e^{\frac{2\pi\mathrm{i}k(z-\overline{a}l)}{b}+\frac{\pi\mathrm{i}((a,b)l+by-az)}{b}}\overline{E}_{j-1}\biggl{(}\frac{(a,b)l+by-az}{b}\biggl{)}$
$\quad+2a^{m}m!n!(a,b)\mathrm{sgn}(a)\sum_{j=1}^{n+1}\binom{m+n+1-j}{m}\frac{b^{n+1-j}}{(j-1)!\bigl{(}2\pi\mathrm{i}(k-\frac{a+b}{2})\bigl{)}^{m+n+2-j}}$
×$\qquad\times\sum_{l=0}^{|\frac{a}{(a,b)}|-1}e^{\frac{2\pi\mathrm{i}k(y+\overline{b}l)}{a}-\frac{\pi\mathrm{i}((a,b)l+by-az)}{a}}\overline{E}_{j-1}\biggl{(}-\frac{(a,b)l+by-az}{a}\biggl{)}.$ (3.8)
Inserting (3.4), (3) and (3) into (3.1), and it then follows that
$e^{\pi\mathrm{i}(ax+bx+y+z)}\overline{E}_{m}(ax+y)\overline{E}_{n}(bx+z)$
$\quad+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)\delta_{\mathbb{Z}}(ax+y)\delta_{\mathbb{Z}}(bx+z)$
[…]
×$\qquad\times\frac{a^{m+1-j}}{(j-1)!\bigl{(}2\pi\mathrm{i}((a,b)k-\frac{a+b}{2})\bigl{)}^{m+n+2-j}}$
×$\qquad\times\sum_{l=0}^{|\frac{b}{(a,b)}|-1}e^{\frac{2\pi\mathrm{i}(a,b)k(z-\overline{a}l)}{b}+\frac{\pi\mathrm{i}((a,b)l+by-az)}{b}}\overline{E}_{j-1}\biggl{(}\frac{(a,b)l+by-az}{b}\biggl{)}$
$\quad+2a^{m}m!n!(a,b)\mathrm{sgn}(a)\sum_{j=1}^{n+1}\binom{m+n+1-j}{m}$
×$\qquad\times\frac{b^{n+1-j}}{(j-1)!\bigl{(}2\pi\mathrm{i}((a,b)k-\frac{a+b}{2})\bigl{)}^{m+n+2-j}}$
×$\qquad\times\sum_{l=0}^{|\frac{a}{(a,b)}|-1}e^{\frac{2\pi\mathrm{i}(a,b)k(y+\overline{b}l)}{a}-\frac{\pi\mathrm{i}((a,b)l+by-az)}{a}}\overline{E}_{j-1}\biggl{(}-\frac{(a,b)l+by-az}{a}\biggl{)}\biggl{)}e^{2\pi\mathrm{i}(a,b)kx}$
$\quad+\frac{(-1)^{n+1}2m!n!(a,b)^{m+n+2}e^{\pi\mathrm{i}(ax+bx+y+z)}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{by-az}{(a,b)}\biggl{)},$
which together with $2(a,b)\mid(a+b)$ and $2\nmid(\overline{a}-\overline{b})$ yields
$\overline{E}_{m}(ax+y)\overline{E}_{n}(bx+z)+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)(-1)^{ax+bx+y+z}\delta_{\mathbb{Z}}(ax+y)\delta_{\mathbb{Z}}(bx+z)$
[…]
×$\qquad\times\frac{a^{m+1-j}}{(j-1)!(a,b)^{m+n+1-j}(2\pi\mathrm{i}k)^{m+n+2-j}}$
×$\qquad\times\sum_{l=0}^{|\frac{b}{(a,b)}|-1}(-1)^{l}e^{\frac{2\pi\mathrm{i}(a,b)k(z-\overline{a}l)}{b}}\overline{E}_{j-1}\biggl{(}\frac{(a,b)l+by-az}{b}\biggl{)}$
$\quad+2a^{m}m!n!\mathrm{sgn}(a)\sum_{j=1}^{n+1}\binom{m+n+1-j}{m}$
×$\qquad\times\frac{b^{n+1-j}}{(j-1)!(a,b)^{m+n+1-j}(2\pi\mathrm{i}k)^{m+n+2-j}}$
×$\qquad\times\sum_{l=0}^{|\frac{a}{(a,b)}|-1}(-1)^{l}e^{\frac{2\pi\mathrm{i}(a,b)k(y+\overline{b}l)}{a}}\overline{E}_{j-1}\biggl{(}-\frac{(a,b)l+by-az}{a}\biggl{)}\biggl{)}e^{2\pi\mathrm{i}(a,b)kx}$
$\quad+\frac{(-1)^{n+1}2m!n!(a,b)^{m+n+2}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{by-az}{(a,b)}\biggl{)}.$ (3.9)
Hence, we see from (2.2) and (3) that
$\overline{E}_{m}(ax+y)\overline{E}_{n}(bx+z)+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)(-1)^{ax+bx+y+z}\delta_{\mathbb{Z}}(ax+y)\delta_{\mathbb{Z}}(bx+z)$
$=-2b^{n}\mathrm{sgn}(b)\sum_{j=1}^{m+1}\binom{m}{j-1}\frac{a^{m+1-j}}{(a,b)^{m+n+1-j}(m+n+2-j)}$
×$\qquad\times\sum_{l=0}^{|\frac{b}{(a,b)}|-1}(-1)^{l}\overline{E}_{j-1}\biggl{(}\frac{(a,b)l+by-az}{b}\biggl{)}$
×$\qquad\times\overline{B}_{m+n+2-j}\biggl{(}(a,b)x+\frac{(a,b)(z-\overline{a}l)}{b}\biggl{)}$
$\quad-2a^{m}\mathrm{sgn}(a)\sum_{j=1}^{n+1}\binom{n}{j-1}\frac{b^{n+1-j}}{(a,b)^{m+n+1-j}(m+n+2-j)}$
×$\qquad\times\sum_{l=0}^{|\frac{a}{(a,b)}|-1}(-1)^{l}\overline{E}_{j-1}\biggl{(}-\frac{(a,b)l+by-az}{a}\biggl{)}$
×$\qquad\times\overline{B}_{m+n+2-j}\biggl{(}(a,b)x+\frac{(a,b)(y+\overline{b}l)}{a}\biggl{)}$
$\quad+\frac{(-1)^{n+1}2m!n!(a,b)^{m+n+2}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{by-az}{(a,b)}\biggl{)}.$ (3.10)
Since $\bigl{(}\frac{a}{(a,b)},\frac{b}{(a,b)}\bigl{)}=1$, $\frac{al}{(a,b)}$ runs over a complete residue system modulo $|\frac{b}{(a,b)}|$ as $l$ does, and $-\frac{bl}{(a,b)}$ runs over a complete residue system modulo $|\frac{a}{(a,b)}|$ as $l$ does. It follows from (3) that
$\overline{E}_{m}(ax+y)\overline{E}_{n}(bx+z)+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)(-1)^{ax+bx+y+z}\delta_{\mathbb{Z}}(ax+y)\delta_{\mathbb{Z}}(bx+z)$
$=-2b^{n}\mathrm{sgn}(b)\sum_{j=1}^{m+1}\binom{m}{j-1}\frac{a^{m+1-j}}{(a,b)^{m+n+1-j}(m+n+2-j)}$
×$\qquad\times\sum_{l=0}^{|\frac{b}{(a,b)}|-1}(-1)^{l}\overline{E}_{j-1}\biggl{(}\frac{a(l-z)}{b}+y\biggl{)}\overline{B}_{m+n+2-j}\biggl{(}(a,b)x-\frac{(a,b)(l-z)}{b}\biggl{)}$
$\quad-2a^{m}\mathrm{sgn}(a)\sum_{j=1}^{n+1}\binom{n}{j-1}\frac{b^{n+1-j}}{(a,b)^{m+n+1-j}(m+n+2-j)}$
×$\qquad\times\sum_{l=0}^{|\frac{a}{(a,b)}|-1}(-1)^{l}\overline{E}_{j-1}\biggl{(}\frac{b(l-y)}{a}+z\biggl{)}\overline{B}_{m+n+2-j}\biggl{(}(a,b)x-\frac{(a,b)(l-y)}{a}\biggl{)}$
$\quad+\frac{(-1)^{n+1}2m!n!(a,b)^{m+n+2}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{by-az}{(a,b)}\biggl{)}.$ (3.11)
It is easily shown from (2.1), Lemma 2.2, the property of residue systems and the division algorithm that
$\frac{1}{(a,b)^{m+n+1-j}}\sum_{l=0}^{|\frac{b}{(a,b)}|-1}(-1)^{l}\overline{E}_{j-1}\biggl{(}\frac{a(l-z)}{b}+y\biggl{)}$
×$\quad\times\overline{B}_{m+n+2-j}\biggl{(}(a,b)x-\frac{(a,b)(l-z)}{b}\biggl{)}$
$=\sum_{k=0}^{(a,b)-1}\sum_{l=0}^{|\frac{b}{(a,b)}|-1}(-1)^{k|\frac{b}{(a,b)}|+l}\overline{E}_{j-1}\biggl{(}\frac{a\bigl{(}k|\frac{b}{(a,b)}|+l-z\bigl{)}}{b}+y\biggl{)}$
×$\quad\times\overline{B}_{m+n+2-j}\biggl{(}x-\frac{l-z+k|\frac{b}{(a,b)}|}{b}\biggl{)}$
$=\sum_{l=0}^{|b|-1}(-1)^{l}\overline{E}_{j-1}\biggl{(}\frac{a(l-z)}{b}+y\biggl{)}\overline{B}_{m+n+2-j}\biggl{(}x-\frac{l-z}{b}\biggl{)}.$ (3.12)
Hence, by applying (3) to the right hand side of (3), we arrive at
$\overline{E}_{m}(ax+y)\overline{E}_{n}(bx+z)$
$=-2b^{n}\mathrm{sgn}(b)\sum_{j=0}^{m}\binom{m}{j}\frac{a^{m-j}}{m+n+1-j}\sum_{l=0}^{|b|-1}(-1)^{l}\overline{E}_{j}\biggl{(}\frac{a(l-z)}{b}+y\biggl{)}$
×$\qquad\times\overline{B}_{m+n+1-j}\biggl{(}x-\frac{l-z}{b}\biggl{)}$
$\quad-2a^{m}\mathrm{sgn}(a)\sum_{j=0}^{n}\binom{n}{j}\frac{b^{n-j}}{m+n+1-j}\sum_{l=0}^{|a|-1}(-1)^{l}\overline{E}_{j}\biggl{(}\frac{b(l-y)}{a}+z\biggl{)}$
×$\qquad\times\overline{B}_{m+n+1-j}\biggl{(}x-\frac{l-y}{a}\biggl{)}$
$\quad+\frac{(-1)^{n+1}2m!n!(a,b)^{m+n+2}}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{by-az}{(a,b)}\biggl{)}$
$\quad-\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)(-1)^{ax+bx+y+z}\delta_{\mathbb{Z}}(ax+y)\delta_{\mathbb{Z}}(bx+z).$ (3.13)
It is clear from (2.2) that for $n\in\mathbb{N}$, $x\in\mathbb{R}$,
$\overline{B}_{n}(-x)=(-1)^{n}\overline{B}_{n}(x).$ (3.14)
Therefore, by substituting $-y$ for $y$, and $-z$ for $z$ in (3), in view of (3.14), we achieve (1.10) and finish the proof of Theorem 1.1.
The proof of Theorem 1.3. It is easily seen from Lemma 2.2 and (3.14) that for $j\in\mathbb{N}$, $c\in\mathbb{Z}\setminus\{0\}$, $x\in\mathbb{R}$,
$\sum_{r=0}^{|c|-1}\overline{B}_{j}\biggl{(}\frac{r+x}{c}\biggl{)}=c^{1-j}\mathrm{sgn}(c)\overline{B}_{j}(x).$ (3.15)
If we replace $x$ by $-(r-x)/c$ and then make the operation $\sum_{r=0}^{|c|-1}$ on both sides of (1.10), then we obtain from (3.15) that
$2b^{n}\mathrm{sgn}(bc)\sum_{j=0}^{m}\binom{m}{j}\frac{(-1)^{j}a^{m-j}}{c^{m+n-j}(m+n+1-j)}S_{j,m+n+1-j}^{(5)}\left(\begin{matrix}a&c&b\\ y&x&z\end{matrix}\right)$
$\quad+2a^{m}\mathrm{sgn}(ac)\sum_{j=0}^{n}\binom{n}{j}\frac{(-1)^{j}b^{n-j}}{c^{m+n-j}(m+n+1-j)}S_{j,m+n+1-j}^{(5)}\left(\begin{matrix}b&c&a\\ z&x&y\end{matrix}\right)$
$=(-1)^{m+n}\sum_{r=0}^{|c|-1}\overline{E}_{m}\biggl{(}-\frac{a(r-x)}{c}-y\biggl{)}\overline{E}_{n}\biggl{(}-\frac{b(r-x)}{c}-z\biggl{)}$
$\quad+\frac{(-1)^{m}2\mathrm{sgn}(c)m!n!(a,b)^{m+n+2}c}{a^{n+1}b^{m+1}(m+n+1)!}\overline{E}_{m+n+1}\biggl{(}\frac{az-by}{(a,b)}\biggl{)}$
$\quad+\delta_{0,m}\delta_{0,n}\mathrm{sgn}(ab)\sum_{r=0}^{|c|-1}(-1)^{-\frac{(a+b)(r-x)}{c}-y-z}$
×$\qquad\times\delta_{\mathbb{Z}}\biggl{(}-\frac{a(r-x)}{c}-y\biggl{)}\delta_{\mathbb{Z}}\biggl{(}-\frac{b(r-x)}{c}-z\biggl{)},$
as desired. This concludes the proof of Theorem 1.3.